# Digital Network Twins for Next-generation Wireless: Creation, Optimization, and Challenges

### v1

@@ I Introduction
To resolve these practical challenges, several surveys and tutorials about DNTs have been proposed [2, 3, 4, 5, 6, 7, 8, 9], mainly describing the deployment of DNT technologies in various service sectors. [10, 3, 4] present a broad survey to explore the integration of DTs into networks for potential application scenarios, such as aviation, healthcare, 6G networks, and intelligent transportation systems. [5] focuses on deploying DNTs at edge to improve performance such as user throughput and security, and reducing the cost of communication, computation, and caching. [6, 8] outline the recent trends in machine learning and explore their integration into DNTs for efficient network management. [7] analyzes the current state of the DNT paradigm and classifies the potential threats associated with it, considering its functionality layers and the operational requirements to ensure the trustworthy use of a DT. Specifically, [9] takes more complex information into account, capturing and representing network dynamics as well as stimulating with the modeled reality in real-time. Nevertheless, despite the widespread interest in deploying DNTs for telecommunication, it has not yet been comprehensively studied about the full lifecycle of DNTs, particularly in fine-grained creation, real-time adaptation, resource deployment optimization, and security.
This article presents a broad overview of DNTs and unveils the full potential of their integration into communication networks, covering the fundamentals, the emergent applications, and critical challenges in multiple dimensions. We start by exploring challenges and opportunities in DNT model creation and adaption, in the aspects of mapping solutions, twinning data generation, and real-time synchronization requirements. Then, we consider the joint optimization of mapping and resource management to deploy the DNTs in practical network systems, by integrating split learning, graph neural networks, and proximal policy optimization within a reinforcement learning framework. Moreover, the security issues within the DNTs are unveiled, with a particular focus on adversarial attacks and information leakage. Lastly, we showcase concrete application scenarios of DNTs supported by use-case experiments. The paper concludes by outlining future research directions in addressing the above technical challenges.
@@ II Preliminaries and Overview
A digital twin (DT) is a virtual representation of a physical system designed to reflect its runtime process and critical properties. It covers the entire lifecycle of each object, continuously updating from real-time data, and employs reasoning to predict future dynamics and inform decision-making. As an extension, DNTs offer advanced capabilities beyond traditional DTs by incorporating network-specific elements. While DTs enable continuous data exchange between isolated digital models and physical systems, DNTs extend this by jointly modeling, simulating, and optimizing complex communication networks in real time. DNTs simulate network behaviors such as traffic patterns, latency, and data transmission efficiency while adapting to dynamic network topologies and ensuring quality of service (QoS). They address multi-dimensional challenges such as inter- and intra-connectivities, security, and resource management more effectively than DTs, which typically focus on individual systems. A key feature of DNTs is their ability to manage transitions between multiple twins within a network and dynamically allocate resources—such as bandwidth and computing power—across interconnected twins. This is critical in large-scale, dynamic networks where traffic, connectivity, and user demands fluctuate. DNTs ensure efficient data flow, resilience, and adaptability across the network, coordinating interactions between multiple twins to optimize performance.
@@
As shown in Fig. 1, a physical network (PN) consists of user devices (i.e., virtual reality headsets, sensors, servers, base stations (BSs), etc), network protocols, and communication channels. Specifically, a user device is defined as a physical network object (PNO) that can be described by a set of internal properties (i.e., physical dimension, location, mobility patterns) and a set of external properties (i.e., data transmission, access channel). In particular, external properties are used to capture how PNO affects other PNOs in the PN. On the other hand, the DNT, consisting a network of twins, is a real-time virtual mapping of the PN such that it is expected to comprise the same components with the same internal and external properties. Specifically, the DNT is built based on the following two processes: 1) Mapping all PNOs with their internal and external properties as well as their interrelations into a digital environment; 2) Self-evolving this digital environment according to the changes in any internal or external properties of PNOs in the PN. For example, a user device switches its data transmission from BS1 to BS2 in the PN, which will also affect the external properties (e.g., spectrum access) of other user devices due to the varying communication interference. Therefore, when the DNT updates the external property for a mobile device, it must also update the properties for other affected user devices accordingly. This considered representation model will guide the fundamental designs of DNT creation, updates, and applications as discussed in the following sections. Notably, while the existing works [10, 5, 6, 7, 8, 9] extensively explored the concept and applications of DNT architectures, such as simulating 5G networks and optimizing spectrum usage, practical challenges regarding efficient creation and real-time updates remain underdeveloped. For instance, a conceptual survey on DNTs in [10] highlights how DTs enhance network performance through real-time monitoring and predictive analytics, emphasizing their role in intelligent network management. Another recent survey [7] explores the security landscape of DTs by categorizing potential threats and vulnerabilities in various downstream applications. It mainly emphasizes the increasing attack surface introduced by DTs, particularly due to their integration with physical and cyber domains. Orthogonal to existing literature, this article specifically addresses the foundational building blocks and online adaptation of DNTs within communication networks, introducing novel mapping mechanisms and resource deployment strategies to address a network of DTs. We also investigate continuous monitoring and optimization to mitigate security and privacy concerns in critical network settings.
Figure 2: Summary of the challenges and opportunities of the DNT.
II-B DNT Applications
II-B1 What-if Analysis and Troubleshooting
DNT offers a unique safe sandbox to simulate various runtime network statuses and configurations, where the impact and potential loss of dangerous operations can be safely evaluated and analyzed. In this context, DNT not only facilitates the training of network operators in countering emergent service disruptions and cybersecurity issues but also serves as a virtual environment for the detection of network misconfiguration, performance bottleneck, and link failures without any impact on the physical network. Notably, proactive measures powered by DNT excel conventional reactive measures for network fault recovery, which often suffers from a longer delay to prevent lasting severe economic loss. For example, the prediction of poor link quality for a mobile device availing URLLC service, at a certain point in the future (time or distance) will enable the network to prepare for such an event with proactive measures such as allocating more resources, switching to different frequency bands, preparing for new beams, etc.
II-B2 Network Planning and Management
The increasing scale and complexity of modern networks pose challenges in meeting the demands of larger networks efficiently. By utilizing DNT, ISPs can predict when existing networks will exhaust their resources given a specified growth in users. Furthermore, these performance estimates offer invaluable insights for planning optimal upgrades capable of handling such growth effectively. In essence, the DNT empowers network operators to make informed decisions and anticipate the need for network upgrades. Notably, a DT prototype developed within the Hexa-X initiative [11] introduces dynamic control of 5G slicing, ensuring that QoS guarantees are upheld for each slice, catering to diverse traffic patterns. It also facilitates resource allocation for each slice in accordance with service level agreements (SLAs). Furthermore, the DNT envisions equitable capacity assignment and real-time planning for future 6G infrastructure [8], underscoring its adaptability and foresight in shaping the landscape of upcoming DT-assisted network technologies.
@@ III Challenges and Opportunities
This section addresses the challenges, mitigation strategies, and prospects of DNTs, as summarized in Fig. 2. We begin by examining the creation and adaptation of DNTs, with a particular emphasis on optimizing operational efficiency. Next, we explore their practical application in wireless resource management, where the continuous monitoring, assessment, and stabilization of DNT quality at scale pose significant challenges. Lastly, we highlight the security risks that emerge when integrating machine learning models into DNTs, especially in privacy-sensitive and mission-critical cyber infrastructures, acknowledging these as critical concerns.
@@ III-A3 Twinning Data Generation
One main challenge of utilizing AI/ML methods to construct DNTs is the necessity to generate a substantial volume of representative data capturing both environmental intricacies and coupled network attributes. However, this task is often impractical and unfeasible in real-world networks due to the extensive measurement campaigns required. Thus, one potential solution is to synthetically generate high-quality data covering a wide range of fine-grained network scenarios, which is then used to develop a faithful, efficient approach for the DNT creation. For instance, a quasi-deterministic (Q-D) ray tracer [12] can be used to generate various instance data within a known scenario covering different access point (AP)/user locations and protocol settings. Such a model can capture the geometrical properties of the channel for each transceiver and generate the profile of delay, path gain, angle of departure (AoD), angle of arrival (AOA), etc, for every path component. However, the data quality from the ray tracer is highly dependent on the abstracted channel model, which needs to be well-designed to acquire an accurate channel impulse response. In this regard, a small amount of measurement data from real-world experimental testbeds will contribute to the data generation process in two aspects: a) help derive the aligned scenario distributions, and b) create a feedback path to tune the related deterministic channel parameters such that the ray tracer can produce a more accurate dataset that matches the actual measurements.
@@ III-B Wireless Resource Management for Digital Network Twins
Different from current resource management work that optimizes temporal and spatial allocations only for the physical network, resource management herein must consider networking resource (i.e., spectrum, transmit power and computational power) management and mapping method selection to optimize the performance of both physical networks and DNTs. In particular, each mapping method requires unique physical network data to generate DNTs, thereby resulting in different data transmission requirements. For example, to map communication attributes between two entities, horizontal mappings require the digital counterpart to learn the network evolution while vertical mappings mainly require environment parameters from the physical counterpart. In this regard, joint optimization of mapping method selection and resource management yields the following unique challenges.
@@ III-B2 Optimized DNT Deployment
Another challenge is to address the optimization of the DNT performance (e.g., efficiency and accuracy of generation process) via determining mapping methods (i.e., horizontal or vertical mappings introduced in Subsection III A) for each network entity and wireless resource (i.e., bandwidth, transmit power, beamforming parameters) for data transmission from physical networks to DNTs. Since mapping method selection and actual wireless resource allocation are always coupling and time-dependent, where the current mapping and resource allocation will explicitly affect future decisions, one may study the use of reinforcement learning (RL) to maximize the joint utility function derived from the quality assessment stage.
However, implementing RL for joint mapping selection and wireless resource management encounters two main challenges. First, representing the dynamics of twining objects and network attributes as a state in an RL framework is non-trivial, as nextG networks may involve numerous object nodes with diverse potential dynamics, which cannot be adequately captured by a limited-size vector. One promising solution is to use a GNN framework with heterogeneous graphs to represent the network dynamics due to its ability to represent different physical networks and DNTs with a unique embedding vector and to analyze the inherent relationships among them. The second challenge lies in how to deal with the continuous action space over the actual network (e.g., with bandwidth and power allocation, beamforming process). This is because standard RL (e.g., Q-learning) that uses value functions to select optimal actions cannot deal with continuous action space. While discretizing these optimization variables will sacrifice the DNT performance, this motivates the use of proximal policy optimization (PPO) that can stabilize and output an optimal action according to an RL state.
@@ III-B3 Large-scale Twin Optimization
The training and implementation of a centralized RL framework for DNT optimization require the collection of all physical network dynamics and DNT states thus probably leading to high communication overhead and implementation cost, particularly in a large-scale nextG wireless network. As a result, the RL model must be extended to a distributed scheme, referred to as DRL, which enables distributed APs and user devices in a physical network to cooperatively determine their mapping methods and resource utilization thus efficiently optimizing the utility function. To this end, one solution is to explore a split learning (SL) approach to decide on how to divide GNNs and RL models across different network entities, and how to leverage that division to share data and perform collaborative RL for the creation of DNTs. Unlike standard SL, which categorizes neural networks based solely on data distribution and learning performance, DNT uses separated neural network models that take into account factors such as computational resources, transmit power, wireless spectrum resources, and the dynamics of each device. Given the divided neural networks, the next step is to investigate different information-sharing procedures among DNTs and analyze their impacts on collaborative RL performance.
@@ III-C2 Information leakage with ML models in DNTs
Given that ML models form a core component of the DNTs, a significant threat model pertains to ML parameter leaks within the network. This involves potential adversaries like VR users and servers excluded from trustworthy training processes. These adversaries, whether active or passive, pose the risk of inferring raw data from both VR users and servers. The consequential risk is the potential leakage of sensitive information, encompassing user details as well as system-level operational details when leveraging these ML models in virtual environments. One potential solution is to design a novel distributed learning framework that enables distributed wireless devices to collaboratively train ML models while preventing the leakage of sensitive data of connected users during the training of ML models for DNT generation and synchronization. This framework combines quantization theory and federated learning, to enable wireless devices to efficiently generate a globally optimal ML model and prevent information leakage without sacrificing ML performance.
@@ III-C3 Privacy risks within DNTs
The creation and use of DNTs raise significant privacy concerns due to the vast amount of data they capture, process, and store. These concerns encompass unauthorized access, misuse of sensitive information, surveillance, and data breaches. DNTs, being intricate and interconnected, risk inadvertently disclosing personal internet usage, preferences, and intellectual property to unauthorized parties. To address these concerns, robust data encryption and access control measures can be implemented to ensure secure and authorized access to the digital environment. Privacy-by-design principles can also mitigate risks by integrating privacy protections into the DNT mapping phase, including data minimization strategies during physical data collection, which balances the twin fidelity and system security. Additionally, employing techniques like decentralized federated learning among multiple split twins [14] can enhance privacy by distributing model updates instead of raw data to a central physical network controller, and/or with differential privacy adding noise to updates to prevent reverse engineering from DNTs to the physical networks.

### v2

@@ I Introduction
To resolve these practical challenges, several surveys and tutorials about DNTs have been proposed [2, 3, 4, 5, 6, 7, 8, 9, 10], mainly describing the deployment of DNT technologies in various service sectors. [11, 3, 4] present a broad survey to explore the integration of DTs into networks for potential application scenarios, such as aviation, healthcare, 6G networks, and intelligent transportation systems. [5] focuses on deploying DNTs at edge to improve performance, such as user throughput and security, and reducing the cost of communication, computation, and caching. [6, 8] outline the recent trends in machine learning and explore their integration into DNTs for efficient network management. [7] analyzes the current state of the DNT paradigm and classifies the potential threats associated with it, considering its functionality layers and the operational requirements to ensure the trustworthy use of a DT. Specifically, [9] takes more complex information into account, capturing and representing network dynamics as well as interacting with the modeled reality in real-time. Nevertheless, despite the widespread interest in deploying DNTs for telecommunication, it has not yet been comprehensively studied about the full lifecycle of DNTs, particularly in fine-grained creation, real-time adaptation, resource deployment optimization, and security.
This article presents a broad overview of DNTs and unveils the full potential of their integration into communication networks, covering the fundamentals, the emergent applications, and critical challenges in multiple dimensions. We start by exploring challenges and opportunities in DNT model creation and adaptation, in the aspects of mapping solutions, twinning data generation, and real-time synchronization requirements. Then, we consider the joint optimization of mapping and resource management to deploy the DNTs in practical network systems, by integrating split learning, graph neural networks, and proximal policy optimization within a reinforcement learning framework. Moreover, the security issues within the DNTs are unveiled, with a particular focus on adversarial attacks and information leakage. Lastly, we showcase concrete application scenarios of DNTs supported by use-case experiments. The paper concludes by outlining future research directions in addressing the above technical challenges.
@@ II Preliminaries and Overview
A digital twin (DT) is a virtual representation of a physical system designed to reflect its runtime process and critical properties. It covers the entire lifecycle of each object, continuously updating from real-time data, and employs reasoning to predict future dynamics and inform decision-making. As an extension, DNTs offer advanced capabilities beyond traditional DTs by incorporating network-specific elements. While DTs enable continuous data exchange between isolated digital models and physical systems, DNTs extend this by jointly modeling, simulating, and optimizing complex communication networks in real time.
As illustrated in Fig. 1, DNTs differ from traditional DTs by integrating both inter- and intra-connectivity information, including geological distance, coverage area overlap, and data distribution similarity. Together with fine-grained network attributes among components within the physical network (PN), such as user traffic behaviors, the connected DNTs provide a comprehensive view of the entire system. In addition to capturing physical connectivity properties and network dynamics, DNTs maintain a synchronized mapping of the protocol stack states—including PHY/MAC layer parameters (e.g., modulation scheme, channel coding, scheduling policies), transport configurations (e.g., congestion control settings), and application-level QoS requirements—alongside resource states such as spectrum allocation, CPU/GPU utilization, and energy availability. This multi-layer state representation enables DNTs to reflect system-level configurations with high fidelity and temporal accuracy. Leveraging this holistic, cross-layer view, the network optimization problem is no longer solved by isolated DTs; instead, DNTs can coordinate state transitions among multiple twins and dynamically orchestrate resource allocation (e.g., bandwidth, computing cycles, storage) across interconnected twins. Such coordination is essential in large-scale, dynamic networks where traffic patterns, connectivity conditions, and service demands fluctuate.
@@
Specifically, a PN consists of user devices (i.e., virtual reality headsets, sensors, servers, base stations (BSs), etc), network protocols, and communication channels. Specifically, a user device is defined as a physical network object (PNO) that can be described by a set of internal properties (i.e., physical dimension, location, mobility patterns) and a set of external properties (i.e., data transmission, access channel). In particular, external properties are used to capture how PNO affects other PNOs in the PN. On the other hand, the DNT, consisting of a network of twins, is a real-time virtual mapping of the PN such that it is expected to comprise the same components with the same internal and external properties. Specifically, the DNT is built based on the following two processes: 1) Mapping all PNOs with their internal and external properties as well as their interrelations into a digital environment; 2) Self-evolving this digital environment according to the changes in any internal or external properties of PNOs in the PN. For example, a user device switches its data transmission from BS1 to BS2 in the PN, which will also affect the external properties (e.g., spectrum access) of other user devices due to the varying communication interference. Therefore, when the DNT updates the external property for a mobile device, it must also update the properties for other affected user devices accordingly. This considered representation model will guide the fundamental designs of DNT creation, updates, and applications as discussed in the following sections. Notably, while the existing works [11, 5, 6, 7, 8, 9] extensively explored the concept and applications of DNT architectures, such as simulating 5G networks and optimizing spectrum usage, practical challenges regarding efficient creation and real-time updates remain underdeveloped. For instance, a conceptual survey on DNTs in [11] highlights how DTs enhance network performance through real-time monitoring and predictive analytics, emphasizing their role in intelligent network management. Another recent survey [7] explores the security landscape of DTs by categorizing potential threats and vulnerabilities in various downstream applications. It mainly emphasizes the increasing attack surface introduced by DTs, particularly due to their integration with physical and cyber domains. Orthogonal to existing literature, this article specifically addresses the foundational building blocks and online adaptation of DNTs within communication networks, introducing novel mapping mechanisms and resource deployment strategies to address a network of DTs.
TABLE I: Summary of the challenges and opportunities of the DNT.
@@ III Challenges and Opportunities
This section addresses the challenges, mitigation strategies, and prospects of DNTs, as summarized in Fig. I. We begin by examining the creation and adaptation of DNTs, with a particular emphasis on optimizing operational efficiency. Next, we explore their practical application in wireless resource management, where the continuous monitoring, assessment, and stabilization of DNT quality at scale pose significant challenges. Lastly, we highlight the security risks that emerge when integrating machine learning models into DNTs, especially in privacy-sensitive and mission-critical cyber infrastructures, acknowledging these as critical concerns.
@@ III-A Creation and Adaptation of Digital Network Twin
Recent studies typically employ machine learning (ML) methods [12], such as LSTM and CNN, to build the initial DNTs. Combined with network simulators, these models create virtual scenarios that closely mirror the PN. However, However, these ML-based mapping process demands substantial communication and computation, as it typically requires offline training over large-scale datasets and potentially high-dimensional feature exchanges between distributed nodes, preventing them from supporting latency-sensitive services and continuous data synchronization.
@@ III-A3 Twinning Data Generation
One main challenge of utilizing AI/ML methods to construct DNTs is the necessity to generate a substantial volume of representative data capturing both environmental intricacies and coupled network attributes. However, this task is often impractical and unfeasible in real-world networks due to the extensive measurement campaigns required. Thus, one potential solution is to synthetically generate high-quality data covering a wide range of fine-grained network scenarios, which is then used to develop a faithful, efficient approach for the DNT creation. For instance, a quasi-deterministic (Q-D) ray tracer [13] can be used to generate various instance data within a known scenario covering different access points (AP)/user locations and protocol settings. Such a model can capture the geometrical properties of the channel for each transceiver and generate the profile of delay, path gain, angle of departure (AoD), angle of arrival (AOA), etc, for every path component. However, the data quality from the ray tracer is highly dependent on the abstracted channel model, which needs to be well-designed to acquire an accurate channel impulse response. In this regard, a small amount of measurement data from real-world experimental testbeds will contribute to the data generation process in two aspects: a) help derive the aligned scenario distributions, and b) create a feedback path to tune the related deterministic channel parameters such that the ray tracer can produce a more accurate dataset that matches the actual measurements.
@@ III-B Wireless Resource Management for Digital Network Twins
Different from current resource management work that optimizes temporal and spatial allocations only for the physical network, resource management herein must consider networking resource (i.e., spectrum, transmit power, and computational power) management and mapping method selection to optimize the performance of both physical networks and DNTs. In particular, each mapping method requires unique physical network data to generate DNTs, thereby resulting in different data transmission requirements. For example, to map communication attributes between two entities, horizontal mappings require the digital counterpart to learn the network evolution, while vertical mappings mainly require environment parameters from the physical counterpart. In this regard, joint optimization of mapping method selection and resource management yields the following unique challenges.
Current research treats DNTs mainly as testbeds for developing and evaluating existing solution space, with little focus on algorithms that optimize the DNTs’ own resource allocation or related services [12]. For instance, reinforcement-learning policies trained to manage spectrum and computing resources in wireless networks do not transfer directly to DNT environments, where optimization objectives and constraints differ significantly due to limited exploration, inter-realm resource coupling, and non-stationary dynamics as the twin and real network co-evolve.
@@ III-B2 Optimized DNT Deployment
Another challenge is to address the optimization of the DNT performance (e.g., efficiency and accuracy of generation process) via determining mapping methods (i.e., horizontal or vertical mappings introduced in Subsection III A) for each network entity and wireless resource (i.e., bandwidth, transmit power, beamforming parameters) for data transmission from physical networks to DNTs. Since mapping method selection and actual wireless resource allocation are always coupled and time-dependent, where the current mapping and resource allocation will explicitly affect future decisions, one may study the use of reinforcement learning (RL) to maximize the joint utility function derived from the quality assessment stage.
However, implementing RL for joint mapping selection and wireless resource management encounters two main challenges. First, representing the dynamics of twining objects and network attributes as a state in an RL framework is non-trivial, as nextG networks may involve numerous object nodes with diverse potential dynamics, which cannot be adequately captured by a limited-size vector. One promising solution is to use a GNN framework with heterogeneous graphs to represent the network dynamics due to its ability to represent different physical networks and DNTs with a unique embedding vector and to analyze the inherent relationships among them. The second challenge lies in how to deal with the continuous action space over the actual network (e.g., with bandwidth and power allocation, beamforming process). This is because standard RL (e.g., Q-learning) that uses value functions to select optimal actions cannot deal with a continuous action space. While discretizing these optimization variables will sacrifice the DNT performance, this motivates the use of proximal policy optimization (PPO), which can stabilize and output an optimal action according to an RL state.
@@ III-B3 Large-scale Twin Optimization
The training and implementation of a centralized RL framework for DNT optimization require the collection of all physical network dynamics and DNT states, thus probably leading to high communication overhead and implementation cost, particularly in a large-scale nextG wireless network. As a result, the RL model must be extended to a distributed scheme, referred to as DRL, which enables distributed APs and user devices in a physical network to cooperatively determine their mapping methods and resource utilization thus efficiently optimizing the utility function. To this end, one solution is to explore a split learning (SL) approach to decide on how to divide GNNs and RL models across different network entities, and how to leverage that division to share data and perform collaborative RL for the creation of DNTs. Unlike standard SL, which categorizes neural networks based solely on data distribution and learning performance, DNT uses separate neural network models that take into account factors such as computational resources, transmit power, wireless spectrum resources, and the dynamics of each device. Given the divided neural networks, the next step is to investigate different information-sharing procedures among DNTs and analyze their impacts on collaborative RL performance.
@@ III-C Security of Inter-Realm Communication with DNTs
Prior work has concentrated on data-centric and hardware threats to DNTs—e.g., data poisoning and physical tampering [7]. Broader risks like privacy risks, model poisoning, and backdoor attacks have received far less attention, but can degrade inter-realm system performance significantly.
@@ III-C2 Information leakage with ML models in DNTs
Given that ML models form a core component of the DNTs, a significant threat model pertains to ML parameter leaks within the network. This involves potential adversaries like VR users and servers excluded from trustworthy training processes. These adversaries, whether active or passive, pose the risk of inferring raw data from both VR users and servers. The consequential risk is the potential leakage of sensitive information, encompassing user details as well as system-level operational details when leveraging these ML models in virtual environments. One potential solution is to design a novel distributed learning framework that enables distributed wireless devices to collaboratively train ML models while preventing the leakage of sensitive data of connected users during the training of ML models for DNT generation and synchronization. This framework combines quantization theory and federated learning to enable wireless devices to efficiently generate a globally optimal ML model and prevent information leakage without sacrificing ML performance.
@@ III-C3 Privacy risks within DNTs
The creation and use of DNTs raise significant privacy concerns due to the vast amount of data they capture, process, and store. These concerns encompass unauthorized access, misuse of sensitive information, surveillance, and data breaches. DNTs, being intricate and interconnected, risk inadvertently disclosing personal internet usage, preferences, and intellectual property to unauthorized parties. To address these concerns, robust data encryption and access control measures can be implemented to ensure secure and authorized access to the digital environment. Privacy-by-design principles can also mitigate risks by integrating privacy protections into the DNT mapping phase, including data minimization strategies during physical data collection, which balances the twin fidelity and system security. Additionally, employing techniques like decentralized federated learning among multiple split twins [15] can enhance privacy by distributing model updates instead of raw data to a central physical network controller, and/or with differential privacy, adding noise to updates to prevent reverse engineering from DNTs to the physical networks.